The hypotenuse of a right triangle is 30 cm. The length of one leg is 6 cm longer than the other leg. Find the length of the shorter leg.
Thank you

Answer by jorel555(1290)

Let L be the shorter leg. Then:
L^2+(L+6)^2=30^2
L^2+L^2+12L+36=900
2L^2+12L-864=0
L^2+6L-432=0
(L+24)(L-18)=0
L=18,-24
The legs are 18 and 24 cms. long!!!!!!!!!!